Find a counterexample, if possible, to these universally quantified statements, where the domain for all variables consists of all real numbers.

a) ∀x(x2 ( x)

b) ∀x(x2 ( 2)

c) ∀x(|x| > 0


1
Expert's answer
2021-10-04T18:32:41-0400

SOLUTION

a) ∀x(x2 ( x)

solution

"x=2,y=1,x\not= \frac{1}1"

b) ∀x(x2 ( 2)

solution

"x=1,y=20, \\20^2-1>100"

c) ∀x(|x| > 0

solution

"x=8,y=4,\\8^2=4^3"